how do i do these using substitution?
1. 2x-y=-6
4x+y=-6
2. 2x-y=3
4x-y=-1
3. 6x-5y=-1
5x-4y=-1

Answer by Alan3354(69443)

how do i do these using substitution?
1. 2x-y=-6
4x+y=-6
Solve for y in either eqn, sub for it in the other eqn.
-----------
2. 2x-y=3
4x-y=-1
Same as #1.
-----------
3. 6x-5y=-1
5x-4y=-1
Solve for either x or y, then sub in the other eqn.